# SimpleGVR: A Simple Baseline for Latent-Cascaded Video Super-Resolution

### v2

@@ Abstract
Latent diffusion models have emerged as a leading paradigm for efficient video generation. However, as user expectations shift toward higher-resolution outputs, relying solely on latent computation becomes inadequate. A promising approach involves decoupling the process into two stages: semantic content generation and detail synthesis. The former employs a computationally intensive base model at lower resolutions, while the latter leverages a lightweight cascaded video super-resolution (VSR) model to achieve high-resolution output. In this work, we focus on studying key design principles for latter cascaded VSR models, which are underexplored currently. First, we propose two degradation strategies to generate training pairs that better mimic the output characteristics of the base model, ensuring alignment between the VSR model and its upstream generator. Second, we provide critical insights into VSR model behavior through systematic analysis of (1) timestep sampling strategies, (2) noise augmentation effects on low-resolution (LR) inputs. These findings directly inform our architectural and training innovations. Finally, we introduce interleaving temporal unit and sparse local attention to achieve efficient training and inference, drastically reducing computational overhead. Extensive experiments demonstrate the superiority of our framework over existing methods, with ablation studies confirming the efficacy of each design choice. Our work establishes a simple yet effective baseline for cascaded video super-resolution generation, offering practical insights to guide future advancements in efficient cascaded synthesis systems.
@@ 1 Introduction
Recent advancements chen2023seine ; fridman2024scenescape ; yin2023nuwa ; voleti2022mcvd ; blattmann2023align ; zhang2023show ; bartal2024lumiere ; jiang2023text2performer ; polyak2410movie ; ma2025step ; kong2024hunyuanvideo ; seawead2025seaweed ; wan2025wan ; yang2024cogvideox ; zheng2024open in diffusion-based text-to-video (T2V) generation have markedly enhanced the visual quality and coherence of synthesized videos. Leading models, such as Hunyuan kong2024hunyuanvideo , CogXVideo yang2024cogvideox and Wan wan2025wan , rely on DiT backbones with full self-attention to fuse spatial, temporal, and textual cues, producing coherent clips with rich detail. However, their computational cost grows quadratically with spatial resolution: directly generating 1080p output in a single stage demands prohibitive computation and incurs long inference times. Given their multi-resolution training paradigm, running these models at lower resolution (e.g., 5122) strikes an optimal balance between visual fidelity and computational efficiency. We define 512p as a resolution whose total pixel area is nearly 5122. Since 512p outputs capture motion and structure well, applying a lightweight cascaded VSR model to enhance them to 1080p offers a promising solution for efficient high-resolution video synthesis with low computational cost and latency.
@@ 4.2.1 Flow-based degradation
Upon inspecting the first-stage T2V outputs (see Fig. 3), we observe that unlike real-world low-quality videos, these video sequences do not exhibit severe degradations such as severe blur, noise, or compression. Instead, they present two distinctive characteristics: 1) localized motion-dependent blur and (2) frame-to-frame color blending distortion (current-frame colors smeared by previous-frame hues). These two phenomena are closely entangled and cannot be effectively replicated using traditional degradation model realbasicvsr . To synthesize color blending distortion, we first estimate the motion field between adjacent frames using Dense Inverse Search (DIS) dis optical flow algorithm, mapping each point $(x,y)$ in the current frame to its position $(x^{\prime},y^{\prime})$ in the previous frame. Based on the optical flow results, we generate motion masks to identify regions with significant movement and create elliptical patterns within these regions. For each elliptical region, we apply gaussian sampling to extract colors from the previous frame, blend them in RGB space, and generate the colors. We then blend these colors into the current frame using a distance-based weight map, where pixels closer to the ellipse center receive stronger effects.
@@ 4.4 Efficient Pseudo-Global Computation
To reduce computational overhead during inference, we further explore replacing the full attention in SimpleGVR with swin attention. However, despite the use of the shifted window mechanism, this modification leads to noticeable reduction in fine details. We hypothesize that this is due to its limited receptive field and lack of dynamic cross-window interactions. Inspired by MoBA lu2025moba and swin attention’s window partitioning strategy, we divide tokens into 2D window units of fixed size. Each unit performs self-attention within itself and dynamically attends to the top-k most relevant window units across the entire video, including future frames, based on computed relevance scores. This mechanism enables effective long-range modeling and detail preservation, even with small window sizes and small top-k values (e.g., 1).

### v3

# SimpleGVR: A Simple Baseline for Latent-Cascaded Generative Video Super-Resolution

@@ Abstract
Latent diffusion models have emerged as a leading paradigm for efficient video generation. However, as user expectations shift toward higher-resolution outputs, relying solely on latent computation becomes inadequate. A promising approach involves decoupling the process into two stages: semantic content generation and detail synthesis. The former employs a computationally intensive base model at lower resolutions, while the latter leverages a lightweight cascaded video super-resolution (VSR) model to achieve high-resolution output. In this work, we focus on studying key design principles for latter cascaded VSR models, which are underexplored currently. First, we propose two degradation strategies to generate traininpairs that better mimic the output characteristics of the base model, ensuring alignment between the VSR model and its upstream generator. Second, we provide critical insights into VSR model behavior through systematic analysis of (1) timestep sampling strategies, (2) noise augmentation effects on low-resolution (LR) inputs. These findings directly inform our architectural and training innovations. Finally, we introduce interleaving temporal unit and sparse local attention to achieve efficient training and inference, drastically reducing computational overhead. Extensive experiments demonstrate the superiority of our framework over existing methods, with ablation studies confirming the efficacy of each design choice. Our work establishes a simple yet effective baseline for cascaded video super-resolution generation, offering practical insights to guide future advancements in efficient cascaded synthesis systems.
@@ 1 Introduction
Recent advancements chen2023seine ; fridman2024scenescape ; yin2023nuwa ; voleti2022mcvd ; blattmann2023align ; zhang2023show ; bartal2024lumiere ; jiang2023text2performer ; polyak2410movie ; ma2025step ; kong2024hunyuanvideo ; seawead2025seaweed ; wan2025wan ; yang2024cogvideox ; zheng2024open in diffusion-based text-to-video (T2V) generation have markedly enhanced the visual quality and coherence of synthesized videos. Leading models, such as Hunyuan kong2024hunyuanvideo , CogXVideo yang2024cogvideox and Wan wan2025wan , rely on DiT backbones with full self‑attention to fuse spatial, temporal, and textual cues, producing coherent clips with rich detail. However, their computational cost grows quadratically with spatial resolution: directly generating 1080p output in a single stage demands prohibitive computation and incurs long inference times. Given their multi‑resolution training paradigm, running these models at lower resolution (e.g., 5122) strikes an optimal balance between visual fidelity and computational efficiency. We define 512p as a resolution whose total pixel area is nearly 5122. Since 512p outputs capture motion and structure well, applying a lightweight cascaded VSR model to enhance them to 1080p offers a promising solution for efficient high-resolution video synthesis with low computational cost and latency.
@@ 4.2.1 Flow-based degradation
Upon inspecting the first-stage T2V outputs (see Fig. 3), we observe that unlike real-world low-quality videos, these video sequences do not exhibit severe degradations such as severe blur, noise, or compression. Instead, they present two distinctive characteristics: 1) localized motion‑dependent blur and (2) frame‑to‑frame color blending distortion (current‑frame colors smeared by previous‑frame hues). These two phenomena are closely entangled and cannot be effectively replicated using traditional degradation model realbasicvsr . To synthesize color blending distortion, we first estimate the motion field between adjacent frames using Dense Inverse Search (DIS) dis optical flow algorithm, mapping each point $(x,y)$ in the current frame to its position $(x^{\prime},y^{\prime})$ in the previous frame. Based on the optical flow results, we generate motion masks to identify regions with significant movement and create elliptical patterns within these regions. For each elliptical region, we apply gaussian sampling to extract colors from the previous frame, blend them in RGB space, and generate the colors. We then blend these colors into the current frame using a distance-based weight map, where pixels closer to the ellipse center receive stronger effects.
@@ 4.4 Efficient Pseudo-Global Computation
To reduce computational overhead during inference, we further explore replacing the full attention in SimpleGVR with swin attention. However, despite the use of the shifted window mechanism, this modification leads to noticeable reduction in fine details. We hypothesize that this is due to its limited receptive field and lack of dynamic cross-window interactions. Inspired by MoBA lu2025moba and swin attention’s window partitioning strategy, we divide tokens into 2D window units of fixed size. Each unit performs self-attention within itself and dynamically attends to the top‑k most relevant window units across the entire video, including future frames, based on computed relevance scores. This mechanism enables effective long-range modeling and detail preservation, even with small window sizes and small top‑k values (e.g., 1).